Sin 4530°


Sin  4530°


Grade:

3 Answers

Rohan Das
11 years ago

sin4530=sin(25pi+30)

=sin30

=1/2

Akash Kumar Dutta
11 years ago

sin( 25pi + 30)=-sin30= -1/2

Sreesritha
4 years ago
Sin(90×50+30)
As the number is 50 the angle will be in 3rd coordinate 
-sin30
-1/2
Hope this answer will help you in solving.
Thanks.